OpenStudy (anonymous):

Help please Wendy is looking over some data regarding the strength of some rope and how that relates to its number of woven strands. The data seem to be exponential. How can Wendy determine if the data are exponential, and how can she convert that to a logarithmic function?

OpenStudy (anonymous):

@Opcode @KingGeorge @abb0t @wolfe8

OpenStudy (anonymous):

Pleas help me If I fail this I will go below my grade.

ganeshie8 (ganeshie8):

say "y" is strength, and "x" is number of woven strands

OpenStudy (anonymous):

k please do this step by step so ic an take notes

OpenStudy (anonymous):

help me ganeshie8 I em bigpapa021201

OpenStudy (anonymous):

please hes helping me first

OpenStudy (anonymous):

pleas leave

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@ganeshie8 please continue

OpenStudy (anonymous):

@ganeshie8 ??

OpenStudy (anonymous):

i will help you

OpenStudy (anonymous):

Well, Yayo, I'd start with an exponential model based upon the most basic exponential function, y = e^x. Then I'd modify this function by inserting a constant coefficient, k. Then I'd take a couple of points (which I hope you have) and substitute the x and y values for each point into this model: y = ke^x. If you can determine a value k that makes the equation true for all of your data points, then lo and behold, the data are exponentially related. Please try it.

OpenStudy (anonymous):

you just coppied an answer thanks bt no thanks I need @ganeshie8

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

For the data to be exponential, the strength needs to increase very rapidly as the number of woven strands increase, and it needs to follow below equation form :- \(y = Ae^x\)

ganeshie8 (ganeshie8):

sorry, my browser crashed...

OpenStudy (anonymous):

ohh its okay okay whats next?

ganeshie8 (ganeshie8):

@bigpapa021201 after this, il come to ur thread :)

OpenStudy (anonymous):

okay thank you

ganeshie8 (ganeshie8):

\(y = Ab^x\) to convert it to logarithmic function, take "log" both sides : \(\log (y) = \log (Ab^x)\) By using product property, log of products equals sum of logs : \(\log (y) = \log (A) + \log (b^x)\)

ganeshie8 (ganeshie8):

this is a logarithm function now ! note that i have changed "e" to "b", just to be more generic

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

we're done

ganeshie8 (ganeshie8):

see if that makes some sense.... let me knw if smthng doesnt :)

OpenStudy (anonymous):

okay I get it so this would be the answer corretc Ill put it the way i understood it

ganeshie8 (ganeshie8):

yup, just replace the initial "e" wid "b"

ganeshie8 (ganeshie8):

the exponential function is : \(y = Ab^x\)

ganeshie8 (ganeshie8):

its not \(y = Ae^x\) okay ?

OpenStudy (anonymous):

yea i get it thanks :)

OpenStudy (anonymous):

I have another question but

ganeshie8 (ganeshie8):

shoot

OpenStudy (anonymous):

@amistre64 do you know if the formula \[y = a b^{x}\] is correct?

OpenStudy (amistre64):

its hard to parse thru a 3 month old question, can you ask your question anew?

OpenStudy (amistre64):

a log function is not exponential ... but that may be a step in getting a log function simply by inverse

OpenStudy (anonymous):

what exactly do you mean by that?

OpenStudy (anonymous):

HELP! @amistre64

OpenStudy (amistre64):

i mean, instead of asking your question in some outdated thread, use the 'ask a question' box to type in your own question. this one is simply to outdated to for me to focus on any train of thought that may have transpired.

OpenStudy (anonymous):

@amistre64 I tagged you in a new post